OpenStudy (anonymous):

How about the cost of buying a burger (say $ 4 each). Call that variable b, Price of a slice of pizza at some joint? Say $ 3 each. Call that variable p. Cost for me to eat burgers and pizza slices? I'll call that C. C = 4b + 3p

OpenStudy (anonymous):

What is the question that you are asking?

OpenStudy (anonymous):

to write a real world situation with the variable expression

OpenStudy (aripotta):

you just did, didn't you?

OpenStudy (anonymous):

@AriPotta can you help me with a question?

OpenStudy (aripotta):

i guess

OpenStudy (anonymous):

no

OpenStudy (aripotta):

i don't know what you need :/ can you post the whole question?

OpenStudy (anonymous):

i have to make my own real world situation with a variable expression to go with it

OpenStudy (aripotta):

did you come up with the pizza and burger thing? or is that just an example?

OpenStudy (anonymous):

i came up with it but i dont know how to right it as a real world situation like this. Katie is moving to a new house. Small boxes hold 15 pounds each and medium boxes hold 25 pounds each. Write an expression for the total weight in the moving van if Katie uses "s" small boxes and "m" medium boxes.

OpenStudy (aripotta):

The cost of a burger, b, is $4. The cost of a slice of pizza, p, is $3. Write an expression for the cost of a meal including burgers and pizza.

OpenStudy (anonymous):

Evaluate the following expression using the given values. Find x - 3y if x = 3 and y = -2

OpenStudy (aripotta):

just substitute. 3 - 3(-2) = ?

OpenStudy (anonymous):

-2

OpenStudy (aripotta):

i believe it's actually 9

OpenStudy (aripotta):

-3(-2) = 6, so 3 + 6 = 9